OpenStudy (anonymous):

You know the mean and standard deviation of a population. You take a sample from this population and compute the 90% CI for the mean. This interval contains values that are within how many standard deviations of the mean?

OpenStudy (anonymous):

0.90 1.49 1.65 1.83

OpenStudy (anonymous):

1.65

OpenStudy (anonymous):

Thank you :) Would it be to much to ask for help on one more?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Ask away

OpenStudy (anonymous):

Which statement below is true about a 95% confidence interval where the poputation mean is contained within the interval of 3.2 to 3.8? -The width of this interval is 3.8. -The sample mean is 1.19. -The margin of error is 0.3. -The CI of 95% uses 1.960.

OpenStudy (anonymous):

I'm pretty sure it's the margin of error one. The first one isn't true, the second is highly unlikely, and the fourth doesn't make sense.

OpenStudy (anonymous):

I was thinking that, when I did this before I guess a width of .6 and I ended up with a 45% so I'm not exactly taking my word for it :P

OpenStudy (anonymous):

How do you find the value for a chi-squared cdf? I have a table to do and I don't remember the steps to finding it

OpenStudy (anonymous):

Chi-square cdf... hmm, is that like finding the probability between two critical chi square values?

OpenStudy (anonymous):

Like \(P(\chi^2_{.05}\le X\le\chi^2_{.95})\) ? How to find the chi square values?

OpenStudy (anonymous):

http://i.imgur.com/uHpwhPd.gif The table above gives data about Magic Sugar cereal. Which is the value used for chi-squared cdf? (1.5179,999,3) (0.4882,999,4) (0.6287,999,4) (0.4882,999,3) Like for this kind of question

OpenStudy (anonymous):

A wild guess on my part, but it looks like the first value is the chi square test statistic, and the third is the degrees of freedom. No idea what 999 could mean, but since it's in all of the options it might not matter. Calculating the chi square involves using the given data. It'll be the sum of the numbers you find in the bottom row.

OpenStudy (anonymous):

I'm getting \(\chi^2=1.5179\), and the degrees of freedom is definitely 3.

OpenStudy (anonymous):

I got undefined, but I trust you more than me right now, for the next one though I'm getting 2 answers, and neither are shown In a survey done about voting and gender, it was found that 50 of 200 males voted in a recent election while 124 of 220 females voted. What is the difference between the proportion of females who vote and the proportion of males who vote? That is, which is a 95% confidence interval for the difference? 0.225 to 0.402 0.095 to 0.282 0.044 to 0.233 0.050 to 0.147

OpenStudy (anonymous):

Let \(\hat{p}_1=\dfrac{50}{200}=.25\) and \(\hat{p}_2=\dfrac{124}{220}=.564\). The 95% CI will be \[(\hat{p}_1-\hat{p}_2)\pm 1.96\sqrt{\frac{\hat{p}_1(1-\hat{p}_1)}{200}+\frac{\hat{p}_2(1-\hat{p}_2)}{220}}\]

OpenStudy (anonymous):

Okay, like before I got (-0.255, -0.402), would it be A? Its just the negatives...

OpenStudy (anonymous):

The question does mention the difference between female and male voters, so it might actually be \(\hat{p}_2-\hat{p}_1\pm\cdots\). Have you tried that?

OpenStudy (anonymous):

Oh! That would flip the confidence interval! Thank you!!! I'm gonna try finishing this up, I will let you know how I do, thank you so much for everything!!!

OpenStudy (anonymous):

Yup, you're welcome

OpenStudy (anonymous):

OMG OMG OMG!!! I got an 85 THANK YOU!!! I only got the ones I did on my own wrong!!!